OpenStudy (anonymous):

Find the equation of the line that contains the point (4, 3) and is perpendicular to the line y = (3/4)x − 3.

OpenStudy (anonymous):

what is the slope of the line given?

OpenStudy (anonymous):

this is the question right off my assignment.

OpenStudy (anonymous):

what is there is what I have to work with

OpenStudy (anonymous):

ok, what is the slope of y = (3/4)x - 3?

OpenStudy (anonymous):

so am I taking the (3/4)*4, o am I needing to flip it?

OpenStudy (anonymous):

the slope is 3/4 y = m x + b m is the slope so what is the negative reciprocal of 3/4?

OpenStudy (anonymous):

4/3

OpenStudy (anonymous):

almost... have to negate it too

OpenStudy (anonymous):

yea sorry

OpenStudy (anonymous):

negative means change the sign, reciprocal means flip it and that is so much mess, they just say "inverse" and inverse gets very important a couple years from now in your math studies

OpenStudy (anonymous):

ok so its -4/3

OpenStudy (anonymous):

anyway, the slope of the perpendicular line is -4/3 so all we need now is the y intercept so here is where you plug your point into y = -4/3 x + k and solve for k

OpenStudy (anonymous):

ok which I get 3=-5 1/3 + k?

OpenStudy (anonymous):

yes, but you have to solve for k

OpenStudy (anonymous):

8 1/3

OpenStudy (anonymous):

yes, but they don't usually come out like that but when they do, 25/3 is better

OpenStudy (anonymous):

so use that k in y = -4/3 x + k and you should have the right answer

OpenStudy (anonymous):

ok thank you so much